### v2

@@ Acknowledgements
The publication was created within the project of the Minister of Science and Higher Education ”Support for the activity of Centers of Excellence established in Poland under Horizon 2020” on the basis of the contract number MEiN/2023/DIR/3796. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857533. This publication is supported by Sano project carried out within the International Research Agendas programme of the Foundation for Polish Science, co-financed by the European Union under the European Regional Development Fund (J.F.-R., B.W., K.C., and K.R.). The research presented in this paper received support from the funds assigned by Polish Ministry of Science and Technology to AGH University (K.J., K.C., and K.R.). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017739 (J.F.-R., K.J., B.W., and K.R.).
Author contributions
J.F.-R. and K.R. conceptualized, designed, and supervised research; J.F.-R., K.J., B.W., and K.R. designed the software package; K.J. and B.W. implemented the software package and the code to visualize the dendrograms; J.F.-R. designed and performed the experiments with the help of K.J. and B.W.; J.F.-R., K.C., and K.R. analyzed the results; J.F.-R. created the figures and wrote the first draft of the manuscript; all authors revised, edited, and approved the final version of the manuscript.
@@
Modularity maximization and community detection in complex networks through recursive and hierarchical annealing in the D-Wave Advantage quantum processing units
Joan Falcó-Roget1*, Kacper Jurek2,3, Barbara Wojtarowicz1,2, Karol Capała1,2, Katarzyna Rycerz1,2,3
1 Computer Vision, Sano - Centre for Computational Medicine Nawojki 11, Krakow, 30-072, Poland
2 Faculty of Computer Science, AGH University of Krakow Mickiewicza 30, Kraków, 30-059, Poland
3 Quantum Computing Lab, Academic Computer Center Cyfronet AGH, Nawojki 11, Krakow, 30-950, Poland
Corresponding author(s). E-mail(s): j.roget@sanoscience.org, kzajac@agh.edu.pl;
These authors contributed equally to this work
@@ 3 Quantum Computing Lab, Academic Computer Center Cyfronet AGH, Nawojki 11, Krakow, 30-950, Poland
For an undirected network, Newman’s quality function for a given community structure consisting of only two modules is written as [25]:
@@
Quantum annealing (QA) works directly in binary variables $x_{i}=\{0,1\}$ instead of spin-like ones. We perform a simple transformation $s_{i}=2x_{i}-1$ for that. With this in mind, the complete quality function is as follows:
@@
which can be optimized via QA to obtain high-quality divisions of a network into two communities [23, 24, 57, 25]. Thus, we base our method on the fact that QA works well in simple QUBO problems without one-hot encoding.
@@
$\begin{split}B^{\text{g}}_{ij}&=B_{ij}-\delta_{ij}\sum_{k\in G}B_{ik}\\ &=B_{ij}-(1-\gamma)g_{i}\delta_{ij},\\ \end{split}$ (S1.7)
@@
Noteworthy, the two matrices are identical if the resolution parameter is equal to 1 [3].
@@
$\begin{split}B^{d,\text{g}}_{ij}&=B_{ij}^{d}-\delta_{ij}\sum_{k\in G}B_{ik}^{d}\\ &=B_{ij}^{d}-(1-\gamma)g_{i}^{in}\delta_{ij},\\ \end{split}$ (S2.7)
@@
Contrary to the directed modularity matrix B, the sum of all elements of the directed generalized modularity matrix Bd,g is equal to zero, that is, $\forall\gamma$,
@@
This is due to the fact that summing the in- and out-degrees is equivalent. Noteworthy, the two matrices are identical if the resolution parameter is equal to 1 [28].
@@
Proceeding similarly to in the undirected case, The modularity of an undirected network of $|C^{k-1}|$ different communities can be decomposed as
$Q(C^{k-1})=\sum_{c\in\tilde{C}^{k-2}}\left[\frac{L_{c}}{m}-\gamma\left(\frac{g^{in}_{c}g^{out}_{c}}{m^{2}}\right)\right]+\frac{L_{k}}{m}-\gamma\left(\frac{g^{in}_{k}g^{out}_{k}}{m^{2}}\right)$
where $\tilde{C}^{k-2}\cup k=C^{k-1}$. Then, the modularity index of the same network after splitting the $k$-th community in two can also be decomposed as
$Q(C^{k})=\sum_{c\in\tilde{C}^{k-2}}\left[\frac{L_{c}}{m}-\gamma\left(\frac{g^{in}_{c}g^{out}_{c}}{m^{2}}\right)\right]+\sum_{r=1}^{2}\left[\frac{L_{k_{r}}}{m}-\gamma\left(\frac{g^{in}_{k_{r}}g^{out}_{k_{r}}}{m^{2}}\right)\right]$
@@
where we have arbitrarily dropped the constant term $\frac{1-\gamma}{2}$, $\delta_{ij}$ is the Kronecker delta function, we used $s_{i}^{2}=1$, and
@@
[…]

### v3

@@ Acknowledgments
The publication was created within the project of the Minister of Science and Higher Education ”Support for the activity of Centers of Excellence established in Poland under Horizon 2020” on the basis of the contract number MEiN/2023/DIR/3796. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857533. This publication is supported by the Sano project carried out within the International Research Agendas programme of the Foundation for Polish Science, co-financed by the European Union under the European Regional Development Fund (J.F.-R., B.W., K.C., and K.R.). The research presented in this paper received support from the funds assigned by the Polish Ministry of Science and Technology to AGH University (K.J., K.C., and K.R.). We gratefully acknowledge the Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within the computational grant no. PLG/2025/018857 (J.F.-R., K.J., B.W., and K.R.), PLG/2024/017319 (J.F.-R.), and PLG/2024/017418 (B.W.).
@@
For an undirected network, Newman’s quality function for a given community structure consisting of only two modules is written as [Wierzbinski2023]:
@@
Quantum annealing (QA) can also work in binary variables $x_{i}=\{0,1\}$ instead of spin-like ones. We perform a simple transformation $s_{i}=2x_{i}-1$ for that. With this in mind, the complete quality function is as follows:
@@
which can be optimized via QA to obtain high-quality divisions of a network into two communities [Ushijima-Mwesigwa2017, Negre2020, Nembrini2022, Wierzbinski2023]. Thus, we base our method on the fact that QA works well in simple QUBO problems without one-hot encoding.
@@
[…]
@@
Noteworthy, the two matrices are identical if the resolution parameter is equal to 1 [Newman2006].
@@
[…]
@@
Contrary to the directed modularity matrix Bd, the sum of all elements of the directed generalized modularity matrix Bd,g is equal to zero, that is, $\forall\gamma$,
@@
This is due to the fact that summing the in- and out-degrees is equivalent. Noteworthy, the two matrices are identical if the resolution parameter is equal to 1 [Leicht2008].
@@
Proceeding similarly to the undirected case, the modularity of an undirected network of $|C^{k-1}|$ different communities can be decomposed as
$Q(C^{k})=\sum_{c\in\tilde{C}^{k-1}}\left[\frac{L_{c}}{m}-\gamma\left(\frac{g^{in}_{c}g^{out}_{c}}{m^{2}}\right)\right]+\frac{L_{k}}{m}-\gamma\left(\frac{g^{in}_{k}g^{out}_{k}}{m^{2}}\right).$
Then, the modularity index of the same network after splitting the $k$-th community in two can also be decomposed as
$Q(C^{k+1})=\sum_{c\in\tilde{C}^{k-1}}\left[\frac{L_{c}}{m}-\gamma\left(\frac{g^{in}_{c}g^{out}_{c}}{m^{2}}\right)\right]+\sum_{r=1}^{2}\left[\frac{L_{k_{r}}}{m}-\gamma\left(\frac{g^{in}_{k_{r}}g^{out}_{k_{r}}}{m^{2}}\right)\right]$
@@
where $\delta_{ij}$ is the Kronecker delta function, we used $s_{i}^{2}=1$, and
@@
is the directed generalized modularity matrix Bdd,k for the $k$-th community [Newman2006]. In the second row, we have applied a linear mapping from the Ising variables to a set of binary ones $s_{i}=2x_{i}-1$ to obtain the QUBO formulation. The last row follows from some of the properties of the generalized modularity matrix outlined in the previous section in the Supplements.